# Pushing Cops and Robber on Graphs of Maximum Degree 4

### v1

@@ 3. Graphs with Maximum Degree 4
In [DGR+23], the authors established that the if the underlying graph of an oriented graph $\overrightarrow{G}$ is subcubic, then ${\sf c_{sp}}(\overrightarrow{G})=1$. In this section, we extend this result to show that if $\overrightarrow{G}$ is an orientation of a graph $G$ with $\Delta(G)\leqslant 4$, then ${\sf c_{sp}}(\overrightarrow{G})=1$. To prove this result, we first establish this result for the orientations of $4$-regular graphs.
3.1. 4-regular graphs
In this section, we show that if $\overrightarrow{G}$ is an orientation of a 4-regular graph $G$, then $\overrightarrow{G}$ is cop-win (i.e., ${\sf c_{sp}}(\overrightarrow{G})=1$). For the purpose of this section, assume that $\overrightarrow{G}$ is an orientation of a 4-regular graph $G$.
A vertex $v\in V(\overrightarrow{G})$ is said to be visited if $v$ was occupied by $\mathcal{R}$ in some previous round of the game. Roughly speaking, $\mathcal{C}$ will use the following strategy to capture $\mathcal{R}$. If $\mathcal{R}$ moves to a vertex that was previously visited by $\mathcal{R}$, then $\mathcal{R}$ will be trapped after at most two more moves of $\mathcal{R}$. Further, to ensure this, $\mathcal{C}$ will maintain the following invariant: for each visited vertex $u$, $|N^{+}(u)|\leqslant 1$. If this invariant is broken, then $\mathcal{R}$ will be trapped after finitely many moves of $\mathcal{R}$. First, we have the following easy lemma.
@@
Let $\mathcal{R}$ occupy a vertex $v$ such that $|N^{+}(v)|=2$ on a cop’s move. If $\mathcal{C}$ pushes one vertex of $N^{+}(v)$, then observe that $|N^{+}(v)|=1$. Now, if $\mathcal{R}$ does not move in the next round to the out-neighbor of $v$, then $\mathcal{C}$ can trap $\mathcal{R}$ using an application of Lemma 3.1. To ease the presentation, for the rest of this section, whenever $\mathcal{R}$ occupies a vertex $v$ such that $|N^{+}(v)|=2$ and $\mathcal{C}$ pushes one vertex of $N^{+}(v)$, we will assume that $\mathcal{R}$ will move to the other out-neighbor of $v$ to avoid getting trapped and we will say that $\mathcal{R}$ is forced to move to the other out-neighbor of $v$.
In the following lemma, we establish that $\mathcal{C}$ can maintain the invariant that for each visited vertex $v$, $|N^{+}(v)|\leqslant 1$ and whenever this invariant will break, $\mathcal{C}$ will trap $\mathcal{R}$ in a finite number of rounds, and hence catch $\mathcal{R}$ after finitely many rounds (due to Proposition 2.1).
@@
Let $\overrightarrow{G}$ be an orientation of a $4$-regular graph $G$. Let $\mathcal{R}$ occupies a vertex $u\in V(\overrightarrow{G})$ on a cop’s move. Then, after the cop’s move, either,
@@
    $\mathcal{C}$ can ensure the invariant that for each visited vertex $w$, $|N^{+}(w)|\leqslant 1$,
@@
We will prove this result using induction. For the base case, we show that if $\mathcal{R}$ begins at the vertex $u$, then the statement of the lemma holds. Let $\mathcal{R}$ begins the game at $u$. At this point, if $|N^{+}(u)|=4$, then $\mathcal{C}$ can trap $\mathcal{R}$ by pushing $u$. If $|N^{+}(u)|=3$, then $\mathcal{C}$ can achieve (1) by pushing $u$. If $|N^{+}(u)|=2$, then let $N^{+}(u)=\left\{x,y\right\}$. Here, $\mathcal{C}$ will push the vertex $x$. If $\mathcal{R}$ moves to $y$, then we achieve (1), and else, $\mathcal{C}$ will push $y$ in the next round, and $\mathcal{R}$ will be trapped at $u$. If $|N^{+}(u)|\leqslant 1$, then $\mathcal{C}$ will trap $\mathcal{R}$ using Lemma 3.1.
Now, let the condition (1) of our lemma is satisfied till $\mathcal{R}$ is at a vertex $v$ (i.e., each visited vertex, including $v$ has out-degree at most one after the cop’s move). Now, let $R$ moves from $v$ to vertex $u\in N^{+}(v)$. Thus, $N^{+}(v)=\left\{u\right\}$. Now, we distinguish the action of $\mathcal{C}$ based on the cardinality of $|N^{+}(u)|$. If $|N^{+}(u)|\leqslant 1$, then $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma 3.1. If $|N^{+}(u)|=3$, then $\mathcal{C}$ will push $u$. Note that, now, $N^{+}(u)=\left\{v\right\}$ and $N^{+}(v)=\emptyset$ (since pushing $u$ can change the orientation of only one arc incident on $v$, i.e., $\overrightarrow{vu}$ is changed to $\overrightarrow{uv}$). Thus, if $\mathcal{R}$ moves to $v$, then $\mathcal{R}$ is automatically trapped at $v$. Otherwise, $\mathcal{C}$ will push $v$ in the next round to trap $\mathcal{R}$ at $u$.
@@
Figure 1. An illustration for Case 1 of 3.4. Once $\mathcal{R}$ moves to $x$, $\mathcal{C}$ traps $\mathcal{R}$ using Lemma 3.1
Case 1. $x$ has at most one out-neighbor other than $y$ (i.e., $|N^{+}(x)|\leqslant 2$): See Figure 1 for an illustration. In this case, $\mathcal{C}$ begins with pushing $y$ when $\mathcal{C}$ moves to $u$. Now, the only vertex $\mathcal{R}$ can move to is $x$ as $N^{+}(u)=\left\{x\right\}$ after $\mathcal{C}$ pushes $y$. If $\mathcal{R}$ does not move in the next round to $x$, then $\mathcal{C}$ traps $\mathcal{R}$ by pushing $x$. Let $\mathcal{R}$ moves to $x$ after $\mathcal{C}$ pushes $y$. Now, $|N^{+}(x)|\leqslant 1$ (since the orientation of $\overrightarrow{xy}$ was reversed to $\overrightarrow{yx}$ when $\mathcal{C}$ pushed $y$). Hence, $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma 3.1.
Case 2. $y$ has at most one out-neighbor (i.e., $|N^{+}(y)|\leqslant 1$): Due to Case 1, we can safely assume that $|N^{+}(x)|=3$, else $\mathcal{C}$ can trap $\mathcal{R}$ using the strategy from Case 1. See Figure 2 for an illustration. $\mathcal{C}$ begins with pushing the vertex $x$, and now $N^{+}(x)=\left\{u\right\}$ and $|N^{-}(x)|=3$. If $\mathcal{R}$ does not move to $y$ in the next round, $\mathcal{C}$ pushes $y$ and traps $\mathcal{R}$ at $u$. Hence, let $\mathcal{R}$ moves to $y$. At this point, either $|N^{+}(y)|=1$ (if $|N^{+}(y)|=0$ when $\mathcal{C}$ pushed $x$) or $|N^{+}(y)|=2$ and $x\in N^{+}(y)$ (if $|N^{+}(y)|=1$ when $\mathcal{C}$ pushed $x$). If $|N^{+}(y)|=1$, then $\mathcal{C}$ will trap $\mathcal{R}$ using Lemma 3.1.
If $|N^{+}(y)|=2$, i.e., $N^{+}(y)=\{w,x\}$, when $\mathcal{R}$ moved to $y$, then $\mathcal{C}$ will use the following strategy to trap $\mathcal{R}$. Now, $\mathcal{C}$ pushes $w$. Again, if $\mathcal{R}$ does not move to $x$ in the next round, then $\mathcal{C}$ will trap $\mathcal{R}$ by pushing $x$. Hence, $\mathcal{R}$ moves to $x$. At this point, consider the neighborhood of $x$. Initially, $|N^{+}(x)|=3$ and $N^{-}(x)=\left\{u\right\}$. Then, we pushed $x$ and we had $|N^{-}(x)|=3$ and $N^{+}(x)=\{u\}$. After this, we pushed the vertex $w$. Now, if $xw$ is not an edge in the underlying graph (i.e., $\overrightarrow{xw}$ was not an arc before we pushed $x$), then we have that $N^{+}(x)=\left\{u\right\}$ when $\mathcal{R}$ moved to $x$, and hence, $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma 3.1. Otherwise, if $xw$ is an edge in the underlying graph (i.e., $\overrightarrow{xw}$ and $\overrightarrow{yw}$ were arcs before we pushed $x$), then we have that $N^{+}(x)=\left\{x,w\right\}$ when $\mathcal{R}$ moves to $x$. Now, we again push $w$. Now, if $\mathcal{R}$ does not move to $u$ in the next round, $\mathcal{C}$ will trap $\mathcal{R}$ by pushing $u$. Hence, $\mathcal{R}$ moves to $u$. Now, observe the neighborhood of $u$ and sequence of push operations performed. Initially (in the beginning of our case), $N^{+}(u)=\left\{x,y\right\}$. Then, $\mathcal{C}$ pushed $x$ once and $w$ twice. Hence, the only effect in the graph is that of pushing $x$, and $N^{+}(u)=\left\{y\right\}$ when $\mathcal{R}$ moves to $u$. Hence, $\mathcal{R}$ will be trapped by an application of Lemma 3.1.
@@
Figure 2. An illustration for Case 2 of 3.4. We illustrate the case when $w\in N^{+}(x)\cap N^{+}(y)$. Figure 2(e) illustrates two steps: $\mathcal{R}$ moving from $y$ to $x$ and then $\mathcal{C}$ pushing $w$. Once $\mathcal{R}$ moves to $u$ from $x$, $\mathcal{C}$ traps $\mathcal{R}$ using Lemma 3.1
Case 3. $|N^{+}(x)|=3$ and $|N^{+}(y)|=2$: In this case, $\mathcal{C}$ begins with pushing $x$. Now, $\mathcal{R}$ has to move to $y$ in the next round, else $\mathcal{C}$ can trap $\mathcal{R}$ by pushing $y$. Notice that when $\mathcal{R}$ moves to $y$, $|N^{+}(y)|=3$ (since the orientation of $\overrightarrow{xy}$ was reversed when we pushed $x$), $N^{-}(y)=\{u\}$, and $N^{+}(u)=\{y\}$. Now, $\mathcal{C}$ pushes $y$ to make $N^{+}(y)=\left\{u\right\}$ and $N^{+}(u)=\emptyset$. If $\mathcal{R}$ does not move to $u$ in the next round, $\mathcal{C}$ will trap $\mathcal{R}$ by pushing $u$, and if $\mathcal{R}$ moves to $u$, then it will be trapped automatically since $N^{+}(u)=\emptyset$.
Finally, to complete the proof of our claim, we observe that our 3 cases are indeed exhaustive when $\overrightarrow{xy}$ is an arc. Since our underlying graph $G$ is 4-regular and $\overrightarrow{ux},\overrightarrow{uy},\overrightarrow{xy}$ are arcs, clearly we have that $|N^{+}(x)|\leqslant 3$ and $|N^{+}(y)|\leqslant 2$. We consider the case $|N^{+}(x)|=3$ and $|N^{+}(y)|=2$ in Case 3, $|N^{+}(x)|\leqslant 2$ in Case 2, and $|N^{+}(x)|\leqslant 1$ in Case 1. This completes our proof. ∎
Hence, due to 3.4, we can assume that $xy$ is not an edge in the underlying graph $G$. Next, we prove an easy but useful claim.
@@
Without loss of generality, let us assume that $|N^{+}(x)|\neq 2$. $\mathcal{C}$ begins with pushing $y$. Notice that it does not change $|N^{+}(x)|$ since $xy\notin E(G)$. Now, if $\mathcal{R}$ does not move to $x$ in the next round, $\mathcal{C}$ can trap $\mathcal{R}$ at $u$ by pushing $x$. Hence, we assume that $\mathcal{R}$ moves to $x$. We distinguish the following two cases:
Case 1. $|N^{+}(x)|\leqslant 1$: $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma 3.1.
Case 2. $|N^{+}(x)=3|$. $\mathcal{C}$ pushes $x$ and now $N^{+}(x)=\left\{u\right\}$. If $\mathcal{R}$ does not move to $u$ in the next round, $\mathcal{C}$ traps $\mathcal{R}$ at $x$ by pushing $u$ and if $\mathcal{R}$ moves to $u$ it gets trapped automatically since $N^{+}(u)=\emptyset$ (since we have pushed both of its out-neighbors). ∎
In the following claim, we establish that if $xy\notin E(G)$, then we can establish the conditions of our lemma. Observe that the following claim, along with 3.4 will complete our proof.
@@
Due to 3.5, unless $|N^{+}(x)|=|N^{+}(y)|=2$, $\mathcal{C}$ can trap $\mathcal{R}$ in at most two moves of the robber. Hence, we assume that $|N^{+}(x)|=|N^{+}(y)|=2$ (and hence none of $x$ or $y$ is visited vertex) for the rest of this proof. Further, if no vertex of $N^{+}(x)$ (resp. $N^{+}(y)$) is a visited vertex, then $\mathcal{C}$ pushes $x$ (resp. $y$), and if $\mathcal{R}$ moves to $y$ (resp. $x$), then observe that we satisfy condition (1) of our claim, and if $\mathcal{R}$ does not move, then $\mathcal{C}$ can trap $\mathcal{R}$ at $u$ by pushing $x$ (resp. $y$). Therefore, for the rest of this proof, we assume that $|N^{+}(x)|=|N^{+}(y)|=2$ and there is at least one vertex in $N^{+}(x)$ that is visited and at least one vertex in $N^{+}(y)$ that is visited. Further, let $N^{+}(x)=\left\{x_{1},x_{2}\right\}$ and $N^{+}(y)=\left\{y_{1},y_{2}\right\}$, and without loss of generality, let us assume that $x_{1}$ and $y_{1}$ are visited vertices. We remark that possibly $\left\{x_{1},x_{2}\right\}\cap\left\{y_{1},y_{2}\right\}\neq\emptyset$.
@@
    Edge $x_{1}x_{2}$ is oriented as $\overrightarrow{x_{1}x_{2}}$ when $\mathcal{R}$ moves to $u$: Observe that, in this case, $N^{+}(x_{1})=\left\{x_{2}\right\}$ when $\mathcal{R}$ moved to $u$ (since $x_{1}$ is a visited vertex and hence, $|N^{+}(x_{1})|\leqslant 1$). Now, $\mathcal{C}$ begins with pushing the vertex $y$. Hence, the robber has to move to $x$ in the next round, as otherwise, $\mathcal{C}$ will push $x$ in the next round to trap $\mathcal{R}$ at $u$. Observe that, at this point, $|N^{+}(x_{1})|\leqslant 2$ as possibly $\overrightarrow{yx_{1}}$ was an arc when $\mathcal{R}$ moved from $v$ to $u$ and then we pushed $y$. Next, $\mathcal{C}$ pushes $x_{2}$, which ensures $|N^{+}(x_{1})|\leqslant 1$ again. Now, if $\mathcal{R}$ does not move to $x_{1}$ in the next round, then $\mathcal{C}$ traps $\mathcal{R}$ by pushing $x_{1}$ and if $\mathcal{R}$ moves to $x_{1}$, it gets trapped via an application of Lemma 3.1 as $|N^{+}(x_{1})|\leqslant 1$.

### v2

@@ 4. Graphs with Maximum Degree 4
In [das2023cops], the authors established that if the underlying graph of an oriented graph $\overrightarrow{G}$ is subcubic, then ${\sf c_{sp}}(\overrightarrow{G})=1$. In this section, we extend this result to show that if $\overrightarrow{G}$ is an orientation of a graph $G$ with $\Delta(G)\leqslant 4$, then ${\sf c_{sp}}(\overrightarrow{G})=1$. Since Lemma˜1.4, along with Proposition˜1.2 and Corollary˜1.3, imply Theorem˜1.5, most of this section is devoted to establish the result for the orientations of $4$-regular graphs (Lemma˜1.4).
A vertex $v\in V(\overrightarrow{G})$ is said to be visited if $v$ was occupied by $\mathcal{R}$ in some previous round of the game. The current position of the robber is also considered visited. Roughly speaking, the cop will apply the following strategy to capture $\mathcal{R}$: the cop ensures that every vertex that was previously visited by the robber is “no longer habitable” for the robber in the sense that if $\mathcal{R}$ moves to a vertex that was previously visited, then $\mathcal{R}$ will be trapped after at most two more moves of $\mathcal{R}$. Thus, the movements of $\mathcal{R}$ in the graph are almost “monotone” except for its last few moves. To ensure this, $\mathcal{C}$ will maintain the following invariant: for each visited vertex $u$, $|N^{+}(u)|\leqslant 1$. If this invariant is broken, then $\mathcal{R}$ will be trapped after finitely many moves of $\mathcal{R}$. First, we have the following easy lemma.
@@
If $\mathcal{R}$ occupies a vertex $v$ such that $|N^{+}(v)|=1$ on the robber’s turn, then $\mathcal{R}$ is forced to move to the unique out-neighbor of $v$, since otherwise, the cop can trap $\mathcal{R}$ by Lemma˜4.1. Whenever $\mathcal{R}$ occupies a vertex $v$ such that $|N^{+}(v)|=2$ and $\mathcal{C}$ pushes one vertex of $N^{+}(v)$, we will assume that $\mathcal{R}$ will move to the other out-neighbor of $v$ to avoid getting trapped, and we will say that $\mathcal{R}$ is forced to move to the other out-neighbor of $v$.
In the following lemma, we establish that $\mathcal{C}$ can maintain the invariant that for each visited vertex $v$, $|N^{+}(v)|\leqslant 1$ and whenever this invariant breaks, $\mathcal{C}$ will trap $\mathcal{R}$ in a finite number of rounds, and hence catch $\mathcal{R}$ after finitely many rounds (due to Proposition˜2.1).
@@
Let $\overrightarrow{G}$ be an orientation of a $4$-regular graph $G$. Let $\mathcal{R}$ occupy a vertex $u\in V(\overrightarrow{G})$ on a cop’s move. Then, after the cop’s move, either,
@@
    the following invariant is satisfied: for each visited vertex $w$, $|N^{+}(w)|\leqslant 1$,
@@
We will prove this result using induction on the number of moves played. For the base case, we show that if $\mathcal{R}$ begins at the vertex $u$, then the statement of the lemma holds. Let $\mathcal{R}$ begin the game at $u$. At this point, if $|N^{+}(u)|=4$, then $\mathcal{C}$ can trap $\mathcal{R}$ by pushing $u$. If $|N^{+}(u)|=3$, then $\mathcal{C}$ can achieve (1) by pushing $u$. If $|N^{+}(u)|=2$, then let $N^{+}(u)=\left\{x,y\right\}$. Here, $\mathcal{C}$ will push the vertex $x$, forcing $\mathcal{R}$ to move to $y$. Notice that (1) is achieved in this case. If $|N^{+}(u)|\leqslant 1$, then $\mathcal{C}$ will trap $\mathcal{R}$ using Lemma 4.1.
Now, suppose the condition (1) of our lemma is satisfied when $\mathcal{R}$ is at a vertex $v$ (i.e., each visited vertex, including $v$ has out-degree at most one after the cop’s move). Now, let $\mathcal{R}$ move from $v$ to some vertex $u\in N^{+}(v)$. Thus, $N^{+}(v)=\left\{u\right\}$. Now, we distinguish the action of $\mathcal{C}$ based on the cardinality of $|N^{+}(u)|$. If $|N^{+}(u)|\leqslant 1$, then $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma 4.1. If $|N^{+}(u)|=3$, then $\mathcal{C}$ will push $u$. Note that, now, $N^{+}(u)=\left\{v\right\}$ and $N^{+}(v)=\emptyset$ (since pushing $u$ can change the orientation of only one arc incident to $v$, i.e., $\overrightarrow{vu}$ is changed to $\overrightarrow{uv}$). Thus, if $\mathcal{R}$ moves to $v$, then $\mathcal{R}$ is automatically trapped at $v$. Otherwise, $\mathcal{C}$ will push $v$ in the next round to trap $\mathcal{R}$ at $u$.
@@
Figure 1. An illustration for Case 1 of ˜$4.$ Once $\mathcal{R}$ moves to $x$, $\mathcal{C}$ traps $\mathcal{R}$ using Lemma˜4.1
Case 1. $x$ has at most one out-neighbor other than $y$ (i.e., $|N^{+}(x)|\leqslant 2$): See Figure 1 for an illustration. In this case, $\mathcal{C}$ begins with pushing $y$ when $\mathcal{C}$ moves to $u$. This forces $\mathcal{R}$ to move to $x$. Now, $|N^{+}(x)|\leqslant 1$ (since the orientation of $\overrightarrow{xy}$ was reversed to $\overrightarrow{yx}$ when $\mathcal{C}$ pushed $y$). Hence, $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma˜4.1.
Case 2. $y$ has at most one out-neighbor (i.e., $|N^{+}(y)|\leqslant 1$): Due to Case 1, we can safely assume that $|N^{+}(x)|=3$, else $\mathcal{C}$ can trap $\mathcal{R}$ using the strategy from Case 1. See Figure˜2 for an illustration. $\mathcal{C}$ begins with pushing the vertex $x$, and now $N^{+}(x)=\left\{u\right\}$ and $|N^{-}(x)|=3$. Thus, $\mathcal{R}$ is forced to move to $y$. At this point, either $|N^{+}(y)|=1$ (if $|N^{+}(y)|=0$ when $\mathcal{C}$ pushed $x$) or $|N^{+}(y)|=2$ and $x\in N^{+}(y)$ (if $|N^{+}(y)|=1$ when $\mathcal{C}$ pushed $x$). If $|N^{+}(y)|=1$, then $\mathcal{C}$ will trap $\mathcal{R}$ using Lemma˜4.1.
If $|N^{+}(y)|=2$, i.e., $N^{+}(y)=\{w,x\}$, when $\mathcal{R}$ moved to $y$, then $\mathcal{C}$ will use the following strategy to trap $\mathcal{R}$. Now, $\mathcal{C}$ pushes $w$, forcing $\mathcal{R}$ to move to $x$. At this point, consider the neighborhood of $x$. Initially, $|N^{+}(x)|=3$ and $N^{-}(x)=\left\{u\right\}$. Then, we pushed $x$ and we had $|N^{-}(x)|=3$ and $N^{+}(x)=\{u\}$. After this, we pushed the vertex $w$. Now, if $xw$ is not an edge in the underlying graph (i.e., $\overrightarrow{xw}$ was not an arc before we pushed $x$), then we have that $N^{+}(x)=\left\{u\right\}$ when $\mathcal{R}$ moved to $x$, and hence, $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma˜4.1. Otherwise, if $xw$ is an edge in the underlying graph (i.e., $\overrightarrow{xw}$ and $\overrightarrow{yw}$ were arcs before we pushed $x$), then we have that $N^{+}(x)=\left\{u,w\right\}$ when $\mathcal{R}$ moves to $x$. At this point, we again push $w$, forcing $\mathcal{R}$ to move to $u$. Now, consider the neighborhood of $u$ and sequence of push operations performed. Initially (in the beginning of our case), $N^{+}(u)=\left\{x,y\right\}$. Then, $\mathcal{C}$ pushed $x$ once and $w$ twice. Hence, the only effect in the graph is that of pushing $x$, and $N^{+}(u)=\left\{y\right\}$ when $\mathcal{R}$ moves to $u$. Hence, $\mathcal{R}$ will be trapped by an application of Lemma˜4.1.
@@
Figure 2. An illustration for Case 2 of ˜$4.$ We illustrate the case when $w\in N^{+}(x)\cap N^{+}(y)$. Figure˜2(e) illustrates two steps: $\mathcal{R}$ moving from $y$ to $x$ and then $\mathcal{C}$ pushing $w$. Once $\mathcal{R}$ moves to $u$ from $x$, $\mathcal{C}$ traps $\mathcal{R}$ using Lemma˜4.1
Case 3. $|N^{+}(x)|=3$ and $|N^{+}(y)|=2$: In this case, $\mathcal{C}$ begins with pushing $x$, forcing $\mathcal{R}$ to move to $y$. Notice that when $\mathcal{R}$ moves to $y$, $|N^{+}(y)|=3$ (since the orientation of $\overrightarrow{xy}$ was reversed when we pushed $x$), $N^{-}(y)=\{u\}$, and $N^{+}(u)=\{y\}$. Now, $\mathcal{C}$ pushes $y$ to make $N^{+}(y)=\left\{u\right\}$ and $N^{+}(u)=\emptyset$, forcing $\mathcal{R}$ to move to $u$, where $\mathcal{R}$ is trapped since $N^{+}(u)=\emptyset$.
Finally, to complete the proof of our claim, we observe that our 3 cases are indeed exhaustive when $\overrightarrow{xy}$ is an arc. Since our underlying graph $G$ is 4-regular and $\overrightarrow{ux},\overrightarrow{uy},\overrightarrow{xy}$ are arcs, clearly we have that $|N^{+}(x)|\leqslant 3$ and $|N^{+}(y)|\leqslant 2$. We consider the case $|N^{+}(x)|=3$ and $|N^{+}(y)|=2$ in Case 3, $|N^{+}(x)|\leqslant 2$ in Case 2, and $|N^{+}(x)|\leqslant 1$ in Case 1. This completes the proof of our claim. ∎
Hence, due to ˜4.4, we can assume that $xy$ is not an edge in the underlying graph $G$. Next, we prove an easy but useful claim.
@@
Without loss of generality, let us assume that $|N^{+}(x)|\neq 2$. $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Notice that this does not change $|N^{+}(x)|$ since $xy\notin E(G)$. We distinguish the following two cases:
Case 1. $|N^{+}(x)|\leqslant 1$: $\mathcal{C}$ can trap $\mathcal{R}$ using Lemma˜4.1.
Case 2. $|N^{+}(x)|=3$. $\mathcal{C}$ pushes $x$, and now $N^{+}(x)=\left\{u\right\}$. This forces $\mathcal{R}$ to move to $u$. Since we have pushed both out-neighbors of $u$, $N^{+}(u)=\emptyset$, and thus $\mathcal{R}$ is trapped. ∎
In the following claim, we establish that if $xy\notin E(G)$, then $\mathcal{C}$ can establish the conditions of our lemma. Observe that the following claim, along with ˜4.4 will complete our proof.
@@
Due to ˜4.5, unless $|N^{+}(x)|=|N^{+}(y)|=2$, $\mathcal{C}$ can trap $\mathcal{R}$ in at most two moves of the robber. Hence, we assume that $|N^{+}(x)|=|N^{+}(y)|=2$ (and hence neither $x$ nor $y$ is a visited vertex) for the rest of this proof. Further, if no vertex of $N^{+}(x)$ (resp. $N^{+}(y)$) is a visited vertex, then $\mathcal{C}$ pushes $x$ (resp. $y$), forcing $\mathcal{R}$ to move to $y$ (resp. $x$), and this satisfies condition (1) of our lemma. Therefore, for the rest of this proof, we assume that $|N^{+}(x)|=|N^{+}(y)|=2$ and that each of $N^{+}(x)$ and $N^{+}(y)$ contains a visited vertex. Further, let $N^{+}(x)=\left\{x_{1},x_{2}\right\}$ and $N^{+}(y)=\left\{y_{1},y_{2}\right\}$, and without loss of generality, let us assume that $x_{1}$ and $y_{1}$ are visited vertices. We remark that possibly $\left\{x_{1},x_{2}\right\}\cap\left\{y_{1},y_{2}\right\}\neq\emptyset$, and we assume that whenever $x_{1}\in\{y_{1},y_{2}\}$, we set $x_{1}=y_{1}$.
@@
    Edge $x_{1}x_{2}$ is oriented as $\overrightarrow{x_{1}x_{2}}$ when $\mathcal{R}$ moves to $u$: Observe that, in this case, $N^{+}(x_{1})=\left\{x_{2}\right\}$ when $\mathcal{R}$ moved to $u$ (since $x_{1}$ is a visited vertex and hence, $|N^{+}(x_{1})|\leqslant 1$). $\mathcal{C}$ begins with pushing the vertex $y$, forcing $\mathcal{R}$ to move to $x$. Observe that, at this point, $|N^{+}(x_{1})|\leqslant 2$ as possibly $\overrightarrow{yx_{1}}$ was an arc when $\mathcal{R}$ moved from $v$ to $u$ and then we pushed $y$. Next, $\mathcal{C}$ pushes $x_{2}$, which ensures $|N^{+}(x_{1})|\leqslant 1$ again. Thus, $\mathcal{R}$ is forced to move to $x_{1}$, where $\mathcal{R}$ will be trapped via an application of Lemma˜4.1 as $|N^{+}(x_{1})|\leqslant 1$.